OpenStudy (anonymous):

The distance from Roy's apartment to the corner store is 1,080 meters. Roy jogged from his apartment to the corner store in 6 minutes. What was Roy's speed, in meters per second?

OpenStudy (tkhunny):

Multiply be the right form of one (1) and you have it. \(\dfrac{1080\;m}{6\;min}\cdot\dfrac{1\;min}{60\;sec}\) q.v. Dimensional Analysis.